OpenStudy (anonymous):

can you do this please: \[\int\limits_{-1}^{0}3x \sqrt{1-3x} dx \] by using the substitution u^2=1-3x

OpenStudy (raden):

u = 1 - 3x ------> x = (1-u)/3 du = -3 dx

OpenStudy (anonymous):

yeah I know that, but then how you can deal with the limits 0 and -1

OpenStudy (raden):

see the integral becomes : 

OpenStudy (raden):

can you evaluate that integration without the limits, first ?

OpenStudy (anonymous):

yeah, yeah.. thanks a lot man

OpenStudy (raden):

welcome :)